# JUICER: Data-Efficient Imitation Learning for Robotic Assembly

### v1

@@ Abstract
While learning from demonstrations is powerful for acquiring visuomotor policies, high-performance imitation without large demonstration datasets remains challenging for tasks requiring precise, long-horizon manipulation. This paper proposes a pipeline for improving imitation learning performance with a small human demonstration budget. We apply our approach to assembly tasks that require precisely grasping, reorienting, and inserting multiple parts over long horizons and multiple task phases. Our pipeline combines expressive policy architectures and various techniques for dataset expansion and simulation-based data augmentation. These help expand dataset support and supervise the model with locally corrective actions near bottleneck regions requiring high precision. We demonstrate our pipeline on four furniture assembly tasks in simulation, enabling a manipulator to assemble up to five parts over nearly 2500 time steps directly from RGB images, outperforming imitation and data augmentation baselines.
@@ I Introduction
In summary, we propose JUICER, a pipeline for learning high-precision manipulation from a small number of demonstrations. Our pipeline combines the benefits of diffusion policy architectures and mechanisms for automatically expanding dataset size via data noising and iterative model development cycles. We show the technique’s effectiveness on four simulated tasks from the FurnitureBench benchmark[1] for long-horizon furniture assembly, including tasks with horizons up to $\sim$2500 timesteps and assemblies of up to 5 parts to be precisely grasped, oriented, and inserted. Our results show that each component of our system improves overall task success compared to vanilla IL baselines. Our ablations illustrate how different design choices impact overall system performance, and we demonstrate that JUICER enables learning high-performance policies using as few as 10 human demonstrations. Finally, we contribute our bottleneck state labeling tool, trajectory augmentation tool, collected demonstration datasets, and IL-based collect-and-infer pipeline for the community to utilize and build on. Supplementary videos and Appendix can be found at https://imitation-juicer.github.io/.
@@ III-C Trajectory Augmentation
Init: $i\leftarrow 1$, $\mathcal{D}_{A}=\varnothing$
while $i<N_{A}$ to $N_{p}$ do
       Sample demonstration trajectory $\tau_{H}\sim\mathcal{D}_{H}$
       Sample bottleneck state $s_{\text{bottleneck}}\sim\tau_{H}$
      Reset world state to $s_{\text{bottleneck}}$
       Sample a state $s_{target}$ within the limits $\mathbf{r},\boldsymbol{\theta},\boldsymbol{\phi}$
       Command EE from $s_{\text{bottleneck}}$ to $s_{target}$ and record inverse actions $\tau_{i}^{\text{cand}}$
       Execute $\tau_{i}^{\text{cand}}$ from $s_{target}$, ending in state $s^{\prime}$
       if $\parallel s^{\prime}-s_{\text{bottleneck}}\parallel^{2}\leq\epsilon$ then
             Store trajectory $\mathcal{D}_{A}\leftarrow\mathcal{D}_{A}\cup\{\tau_{i}^{\text{cand}}\}$
             $i\leftarrow i+1$
@@ IV-A4 Baselines
Figure 4: Average and maximum success rates (%) of methods across tasks. Bolded methods are components of our JUICER pipeline.
Figure 5: Learning one_leg from 10 human demos.
@@ VI-A1 Training Hyperparameters
Parameter Value
Control Mode Position
Action Space Dimension 10
Observation Space Dimension 16
Orientation Representation 6D
Max LR $10^{-4}$
LR Scheduler Cosine
Weight Decay $10^{-6}$
Warmup steps 500
Batch Size 256
Max Epochs 400
Steps per Epoch 400
Image Size Input $2\times 320\times 240\times 3$
Image Size Encoder $2\times 224\times 224\times 3$
Vision Encoder Model ResNet18
Encoder Weights ImageNet 1k
Encoder Parameters $2\times 11$ million
Runs per Condition 5
Encoder Feature Projection Dim 128
@@
Parameter Value
U-Net Down dims [256, 512, 1024]
U-Net Parameters 69 million
Policy total parameters 91 million
Observation Horizon $T_{o}$ 1
Prediction Horizon $T_{p}$ 32
Action Horizon $T_{a}$ 8
DDPM Training Steps 100
DDIM Inference Steps 8
@@
Parameter Value
Residual Blocks 5
Residual Block Width 1024
Parameters 10 million
Policy total parameters 32 million
Observation Horizon $T_{o}$ 1
Prediction Horizon $T_{p}$ (Chunked) 32
Action Horizon $T_{a}$ (Chunked) 8
Prediction Horizon $T_{p}$ (No Chunking) 1
Action Horizon $T_{a}$ (No Chunking) 1

### v2

@@ Abstract
While learning from demonstrations is powerful for acquiring visuomotor policies, high-performance imitation without large demonstration datasets remains challenging for tasks requiring precise, long-horizon manipulation. This paper proposes a pipeline for improving imitation learning performance with a small human demonstration budget. We apply our approach to assembly tasks that require precisely grasping, reorienting, and inserting multiple parts over long horizons and multiple task phases. Our pipeline combines expressive policy architectures and various techniques for dataset expansion and simulation-based data augmentation. These help expand dataset support and supervise the model with locally corrective actions near bottleneck regions requiring high precision. We demonstrate our pipeline on four furniture assembly tasks in simulation, enabling a manipulator to assemble up to five parts over nearly 2500 time steps directly from RGB images, outperforming imitation and data augmentation baselines. Project website: https://imitation-juicer.github.io/.
@@ I Introduction
In summary, we propose JUICER, a pipeline for learning high-precision manipulation from a small number of demonstrations. Our pipeline combines the benefits of diffusion policy architectures and mechanisms for automatically expanding dataset size via data noising and iterative model development cycles. We show the technique’s effectiveness on four simulated tasks from the FurnitureBench benchmark[1] for long-horizon furniture assembly, including tasks with horizons up to $\sim$2500 timesteps and assemblies of up to 5 parts to be precisely grasped, oriented, and inserted. Our results show that each component of our system improves overall task success compared to vanilla IL baselines. Our ablations illustrate how different design choices impact overall system performance, and we demonstrate that JUICER enables learning high-performance policies using as few as 10 human demonstrations. Finally, we contribute our bottleneck state labeling tool, trajectory augmentation tool, collected demonstration datasets, and IL-based collect-and-infer pipeline for the community to utilize and build on.
@@ III-C Trajectory Augmentation
Init: $i\leftarrow 1$, $\mathcal{D}_{A}=\varnothing$ while $i<N_{A}$ to $N_{p}$ do
       Sample demonstration trajectory $\tau_{H}\sim\mathcal{D}_{H}$ Sample bottleneck state $s_{\text{bottleneck}}\sim\tau_{H}$ Reset world state to $s_{\text{bottleneck}}$ Sample a state $s_{target}$ within the limits $\mathbf{r},\boldsymbol{\theta},\boldsymbol{\phi}$ Command EE from $s_{\text{bottleneck}}$ to $s_{target}$ and record inverse actions $\tau_{i}^{\text{cand}}$ Execute $\tau_{i}^{\text{cand}}$ from $s_{target}$, ending in state $s^{\prime}$ if $\parallel s^{\prime}-s_{\text{bottleneck}}\parallel^{2}\leq\epsilon$ then
             Store trajectory $\mathcal{D}_{A}\leftarrow\mathcal{D}_{A}\cup\{\tau_{i}^{\text{cand}}\}$ $i\leftarrow i+1$
@@ IV-A4 Baselines
[Uncaptioned image] Figure 4: Average and maximum success rates (%) of methods across tasks. Bolded methods are components of our JUICER pipeline.
[Uncaptioned image] Figure 5: Learning one_leg from 10 human demos.
@@ VI-A1 Training Hyperparameters
Parameter
Value
Control Mode
Position
Action Space Dimension
10
Observation Space Dimension
16
Orientation Representation
6D
Max LR
$10^{-4}$
LR Scheduler
Cosine
Weight Decay
$10^{-6}$
Warmup steps
500
Batch Size
256
Max Epochs
400
Steps per Epoch
400
Image Size Input
$2\times 320\times 240\times 3$
Image Size Encoder
$2\times 224\times 224\times 3$
Vision Encoder Model
ResNet18
Encoder Weights
ImageNet 1k
Encoder Parameters
$2\times 11$ million
Runs per Condition
5
Encoder Feature Projection Dim
128
@@
Parameter
Value
U-Net Down dims
[256, 512, 1024]
U-Net Parameters
69 million
Policy total parameters
91 million
Observation Horizon $T_{o}$
1
Prediction Horizon $T_{p}$
32
Action Horizon $T_{a}$
8
DDPM Training Steps
100
DDIM Inference Steps
8
@@
Parameter
Value
Residual Blocks
5
Residual Block Width
1024
Parameters
10 million
Policy total parameters
32 million
Observation Horizon $T_{o}$
1
Prediction Horizon $T_{p}$ (Chunked)
32
Action Horizon $T_{a}$ (Chunked)
8
Prediction Horizon $T_{p}$ (No Chunking)
1
Action Horizon $T_{a}$ (No Chunking)
1